if 3ab=6 what is the value of A in terms of B ? : if 3ab=6 what is the value of A in terms of B ?
Answer by orca(336)

We need to solve the equation 3ab = 6 for a.
Dividing both sides of the equation by 3b, we have:
(3ab)/(3b)=6/(3b)
So
a = 2/b